OpenStudy (anonymous):

help? In a certain county, the number of charter schools is 4 less than twice the number of alternative schools. We know that there are 48 charter schools in the county. How many alternative schools are in the county? 52 22 36 26 i tried 48 + 48 = 94 then going to 94 - 4 to get 90 but that's not an answer...???

Vocaloid (vocaloid):

let's define our variables let a = number of alternative schools let c = number of charter schools with me so far?

OpenStudy (anonymous):

yes i am

Vocaloid (vocaloid):

ok, it says that ""number of charter schools is 4 less than twice the number of alternative schools." so c = 2a - 4, with me so far?

OpenStudy (anonymous):

yep. thats what ive already done though.

Vocaloid (vocaloid):

ok, it also says that the number of charter schools is 48 so we have: c = 2a - 4 48 = 2a - 4 do you know how to solve for a now?

OpenStudy (anonymous):

uh- sorry you lost me, 2a would be 2*48 right? and that makes 94? but then take away 4 to get 90? but there is not an option for that

Vocaloid (vocaloid):

noope, not quite

Vocaloid (vocaloid):

we don't know what a is yet, that's what we want to find

Vocaloid (vocaloid):

our equation is 48 = 2a - 4 start by adding 4 to each side.

OpenStudy (anonymous):

ooohh okay so adding five 52=2a-8?

Vocaloid (vocaloid):

nope, not quite.

Vocaloid (vocaloid):

2a-4+4 = ?

OpenStudy (anonymous):

4+4= 8 2a-8?

Vocaloid (vocaloid):

nope -4+4 = ?

OpenStudy (anonymous):

ooooh crud -4 sorry so 0 (im sorry im not the smartest ;-;)

Vocaloid (vocaloid):

don't apologize, you're doing fine :) so 2a-4+4 = 2a now we have 2a = 52 divide both sides by 2 to find a

OpenStudy (anonymous):

26?

Vocaloid (vocaloid):

good! that's our answer ~

OpenStudy (anonymous):

oh that makes a lot more sense then what i was doing thank you so much